# Future-Guided Learning: A Predictive Approach To Enhance Time-Series Forecasting

### v2

@@ 4.1 Regression Forecasting Implementation
Similar to event prediction, the teacher model is pretrained first, followed by inference within the student training loop, where the loss defined in Equation (1) is computed. During model testing, we select the neuron with the highest probability and compute the corresponding MSE, aligning this approach with traditional regression evaluation methods.
We compare two scenarios: $\alpha=0.5$ and $\alpha=0.0$, to assess the effect of using only the teacher’s uncertainty versus combining it with the ground truth labels. For our experiments, we employed a single-layer recurrent neural network (RNN), used a batch size of 1, and trained the teacher model for 20 epochs, while both the student and baseline models were trained for 15 epochs.
Figure 3: Mackey Glass forecasting results with (A) 25 and (B) 50 bins. Results show the MSE loss at each horizon. $\alpha=0$ indicates where the student is trained using the distilled label only, and $\alpha=0.5$ is where a mix of both the distilled and ground truth label are used. A table of results is available in Appendix C.
Figure 4: (a) Discretization Process of Mackey Glass. After data generation, each target value $X_{t+n}$ is assigned to a discrete value based on the number of bins and the range of data. Moreover, the number of output neurons matches the number of bins, such that each bin is treated as a separate class label. (b) Teacher model performs next-step prediction, (c) Student model with FGL forecasts $n=5$ steps into the future, (d) Baseline student model prediction without FGL, also forecasting $5$ steps into the future. FGL allows the student model to capture neighboring spatial information of the target, leading to lower MSE loss and richer representation.
4.2 Regression Forecasting Results
The results show that FGL consistently outperforms the baseline models, achieving an average reduction of 46.4% in MSE with 25 bins, and 48.7% reduction with 50 bins. This improvement is particularly pronounced at $\alpha=0$, where the ground truth labels are completely excluded, and training relies entirely on knowledge distillation from the teacher model. As the forecasting horizon extends, it is expected that the MSE for all cases will increase, but we found that the rate of increase of FGL is slower than that of the baseline. This indicates that FGL exhibits a stronger tolerance to a varying forecasting horizon.
When $\alpha=0$, FGL exhibits the slowest increase in MSE across forecasting horizons. This behavior can be attributed to the student model’s strong reliance on the teacher’s future-oriented knowledge, which helps stabilize the predictions over longer time frames. In contrast, at $\alpha=0.5$, the MSE increases faster than at $\alpha=0$, but still grows at a slower rate compared to the baseline. Here, the partial use of ground truth labels introduces some variability, but the distillation from the teacher continues to provide a stabilizing effect. Eventually, the MSE at $\alpha=0.5$ approaches that of the baseline, indicating a saturation point where the benefits of distillation diminish. Critically, FGL improves performance over the baseline in almost all cases.
4.3 Regression Forecasting Discussion
When compared to the baseline method, FGL improved MSE loss across all forecasting horizons, with the most pronounced gains observed at longer horizons. This trend can be attributed to the nature of the forecasting task: shorter-term predictions inherently have lower error due to their temporal proximity, leaving less room for improvement. In contrast, long-term forecasting is more challenging due to increased uncertainty and error accumulation. Here, the information transferred from the future-oriented teacher model plays a greater role in guiding the student model, helping it make more accurate predictions and mitigate the error growth typical at extended horizons.
A limitation in our regression forecasting formulation is the discretization process (Fig. 4), which can reduce the precision of the model’s predictions. While discretization aids in estimating uncertainty, it also imposes artificial boundaries on the prediction space, potentially limiting the model’s ability to make fine-grained forecasts. To better align with traditional time-series forecasting methods, future work should investigate alternative approaches for uncertainty quantification that avoid the need for discretization.
Figure 5: (A) Illustration of predictive coding in the brain. Information is received by the primary visual cortex (V1), which then propagates to different areas of the brain with more complex levels of abstraction. This propagation takes the form as a predictive estimator: higher level areas aid in prediction of lower level areas. The difference between the prediction and true output is the predictive error. Areas with lighter colors represent low-level abstractions, where darker colors represent the increasing representational demand (thoughts, movement, etc). (B) Hierarchical FGL propagates information via uncertainty, in other words, FGL substitutes the predictive estimator with the uncertainty of each layer. This uncertainty is conveyed via the KL divergence of between each layer’s probability distribution with the successive layer. The difference between each layer’s prediction and true output is the predictive error. In both models, areas lower in the hierarchy process information in a delayed manner, as they are the last to receive it. As the demand for complexity increases, so does the predictive error.
5 Future-Guided Learning And Predictive Coding
The teacher and student models can be described more precisely using a Bayesian prediction framework. Consider a network $f(x_{t+n}|x_{t+n-1},\ldots,x_{1})$, whose objective is to predict the next step $x_{t+n}$ given the preceding sequence $x_{t+n-1},\ldots,x_{1}$. The state of the network up to this point can be expressed as the product of the likelihood $f(x|\theta)$ and the current posterior distribution $\pi(\theta|x_{1},\ldots,x_{t+n-1})$. In essence, the network generates predictions by integrating information accumulated up to the current time step.
${\rm teacher}=f(x_{t+n}|x_{t+n-1},x_{t+n-2},\ldots,x_{1})=\sum f(x_{t+n}|\theta)\cdot\pi(\theta|x_{1},\ldots,x_{t+n-1})$ (2)
${\rm student}=f(x_{t+n}|x_{t},x_{t-1},\ldots,x_{1})=\sum f(x_{t}|\theta)\cdot\pi(\theta|x_{1},\ldots,x_{t})$ (3)
The teacher model has access to data beyond the current time step, leveraging $x_{t+n-1},\ldots,x_{1}$, while the student model is constrained to information available up to $x_{t}$. The divergence between the teacher and student models can be quantified using the KL divergence, which measures how the student’s predictions deviate from the teacher’s more informed predictions:
$D_{KL}({\rm teacher\parallel student})=\sum f(x_{t+n}|x_{t+n-1},\ldots,x_{1})\cdot\log(\frac{f(x_{t+n}|x_{t+n-1},\ldots,x_{1})}{f(x_{t+n}|x_{t},\ldots,x_{1})})$ (4)
The intuition behind the KL divergence lies in quantifying the difference between the predictive distributions of the teacher and student models. When forecasting future events, the student model is inherently less informed than the teacher model, as the teacher has access to an additional $n-1$ time points beyond the student’s observational window. This additional information enables the teacher to construct a more accurate predictive distribution. The KL divergence, therefore, represents the “well-informedness” of the teacher model relative to the student model. As the forecasting horizon increases, the representational complexity for the student grows, as it must account for higher levels of uncertainty and potential chaos in the data.
To address this challenge, we extend the student-teacher framework across a temporal hierarchy, linking models operating at different time steps. The topmost model, associated with the current time step $x_{n+t-1}$, acts as a teacher to the model at $x_{n+t-2}$, which in turn is another teacher to the model at $x_{n+t-3}$, and so on. This relationship propagates downwards, with each model at time step $x_{n+t-k}$ serving as a teacher to the subsequent model at $x_{n+t-(k+1)}$, ultimately connecting to the bottommost model at $x_{t}$. Instead of distilling knowledge directly from the topmost model to the bottommost model, we propose a chained uncertainty propagation through these intermediate models. Each model in the hierarchy operates on a distinct timescale, resulting in temporally stratified representations that progressively refine the forecasting process at varying levels of granularity.

### v3

@@ Methods
Similar to event prediction, the teacher model is pretrained first, followed by inference within the student training loop, where the loss defined in Equation (1) is computed. During model testing, we select the neuron with the highest probability and compute the corresponding MSE, aligning this approach with traditional regression evaluation methods. Full experimental details can be found in Supplementary Note 2.
Future Guided Learning And Predictive Coding
The teacher and student models can be described more precisely using a Bayesian prediction framework. The teacher, with access to $n$ future points, has predictive density
$p_{T}(x_{t+n})\;=\;\int f(x_{t+n}\mid\theta)\,\pi(\theta\mid x_{1:t+n-1})\,d\theta,$
while the student, limited to the current window, uses
$p_{S}(x_{t+n})\;=\;\int f(x_{t+n}\mid\theta)\,\pi(\theta\mid x_{1:t})\,d\theta.$
Their divergence is
$D_{\rm KL}(p_{T}\!\parallel p_{S})=\int p_{T}(x)\,\ln\frac{p_{T}(x)}{p_{S}(x)}\,dx,$
quantifying how much “extra” information the teacher holds over the student as the horizon grows.
Rather than distilling directly from the longest‐horizon teacher, we chain predictions through intermediate models at steps $t+1,\dots,t+n-1$. Each acts as both student (to its predecessor) and teacher (to its successor), propagating uncertainty down the hierarchy.